OpenStudy (babyslapmafro):

Please help me locate all absolute maxima and minima of the following functions: f(x,y)=(x-2)^2+(y+1)^2

OpenStudy (babyslapmafro):

I can see that the minima occurs at (2,-1), but how do I prove this mathematically?

OpenStudy (babyslapmafro):

fx(x,y)=2(x-2) 2(x-2)=0 x=2

OpenStudy (babyslapmafro):

fy(x,y)=2(y+1) 2(y+1)=0 y=-1

OpenStudy (babyslapmafro):

Is that right? how do I prove it's the minimum?

OpenStudy (chillout):

It is right, indeed. Stationary points occur when the first derivative is equal to 0.

OpenStudy (chillout):

fx(x,y)=fy(x,y)=0

OpenStudy (babyslapmafro):

Ok and how do I know it is a min and not a max?

OpenStudy (babyslapmafro):

I know it is a min by looking at the formula but how do I prove this?

OpenStudy (chillout):

The second test derivative. Remember for single variable calculus? It is a maximum if f'(a)=0 and f''(a)<0. For minimum, f''(a)>0.

OpenStudy (chillout):

woops. The second derivative test, I meant.

OpenStudy (babyslapmafro):

what if the second derivative has a variable in it? do I plug the x value I found when fx=0 and then plug that into the second derivative?

OpenStudy (chillout):

It's a bit more complicated than that... But in this case you won't have a variable on it, as fxx is equal to 2, and fyy is 2 as well.

OpenStudy (chillout):

You can have saddle points, which are stationary points but are not minimum nor maximum points.

OpenStudy (chillout):

One second... I have something that might be of your interest.

OpenStudy (chillout):

http://i.imgur.com/ze2jc1m.png

OpenStudy (chillout):

This is the general guideline to check for stationary points.

OpenStudy (babyslapmafro):

ok thanks a lot

OpenStudy (chillout):

No problem. I'm studying Calculus II as well :P